A sample of sulfur vapor weighing 24.4 mg exerts pressure of 135 mm Hg at 293 °C in a 24.9 mL glass bulb. Use this information to determine the molar mass of this sulfur compound.
1
Expert's answer
2020-04-10T13:36:54-0400

Applying formula "pV=nRT"

"\frac{135}{760}\times\frac{24.9}{1000}=\frac{24.4\times10^{-3}}{M}\times0.082\times566\\M=256g"